# LS-EEND: Long-Form Streaming End-to-End Neural Diarization with Online Attractor Extraction

### v1

@@ IV-A2 Real-world Recordings
TABLE II: Real-world Recordings. Utterance (Utt) duration is presented in the format of minimum/average/maximum duration. Ovl stands for overlap ratio.
Dataset Split #Spk #Utt Ovl (%) Duration (mins)
CALLHOME [44] Part 1 2--7 249 17.0 0.86/2.10/10.11
Part 2 2--6 250 16.7 0.77/2.05/10.01
DIHARD II [45] Dev 1--10 192 9.8 0.45/7.44/11.62
Test 1--9 194 8.9 0.63/6.96/13.50
DIHARD III [46] Dev 1--10 254 10.7 0.45/8.07/11.62
Test 1--9 259 9.4 0.63/7.65/13.50
AMI headset mix [47] Train 3--5 136 13.4 7.97/35.59/90.25
@@ IV-B Model Configuration
Following the feature extraction method presented in [12, 26], the input audio features are set to 23-dimensional LogMels for every 10 ms, which are then concatenated with left and right 7 frames, and downsampled by a factor of 10, resulting in a 345-dimensional feature for every 0.1 s. The number of stacked blocks in encoder and decoder are set to $N$=4 and $M$=2, respectively. For both the encoder and decoder, the number of attention heads and hidden units are set to 4 and 256, respectively. The feed-forward dimension is set to 1024 in the encoder and 2048 in the decoder. The convolution kernel size in the Conformer module is set to 16, with a zero-padding of 15 to ensure causality. In the look-ahead module, the kernel size and padding are set to 19 and 9, respectively. Consequently, each frame attends to its previous and future 9 frames, resulting in a latency of 1 s. As already mentioned, the decay factor $\gamma$ in Retention is set to 1.
@@ IV-C Training Setup
The training process is divided into two stages: short-utterance pre-training and long-utterance adaptation. Pre-training is progressively conducted with linearly increased number of speakers, using the very large simulated dataset, and the length of training utterances is set to 100 s. Specifically, the model is trained in multiple steps with the subsets of Sim2spk, Sim{1-4}spk, Sim{1-6}spk and Sim{1-8}spk for 100, 50, 50 and 25 epochs, respectively. Training utterances are randomly extracted from the original utterances in an on-the-fly manner. For each step, the Noam optimizer [41] with 100,000 warm-up steps is utilized.
The pre-trained model cannot directly generalize to longer audio streams and real-world data. At the adaption stage, the pre-trained models are fine-tuned and adapted to specific applications/datasets. The fine-tuning process is also progressively conducted, but now with doubly increased audio lengths, using respective adaption data of each dataset. For real-world datasets, adaption is completed simultaneously for longer audio streams and real-world data. Audio length is increased from 200 s until it approximates the average utterance duration of each dataset. Specifically, audio length is progressively set to 200 s and 400 s for the simulated dataset, to 200 s for CALLHOME, to 200 s and 400 s for DIHARD II and DIHARD III, to 200 s, 400 s, 800 s, and 1600 s for AMI. The adaption utterances are also randomly extracted from the original utterances in an on-the-fly manner. When training with long sequence, the chunk-wise recurrent paradigm described in Section III-B is used for the Retention modules, with a recurrent chunk length of 50 s. The Adam optimizer [48] with a learning rate of $1\times 10^{-5}$ is used. In our experiments, we find that the model fine-tuned using the BCE loss (according to the speaker appearance order) does not converge well when adapted to real-world recordings. Therefore, we utilize the PIT loss [15] for real-world datasets.
@@ IV-D Evaluation Setup
Diarization error rate (DER) is adopted as the evaluation metric. The decision of diarization results is made using a threshold of 0.5. For the simulated data and CALLHOME, a collar tolerance of 0.25 s is used. For DIHARD II, DIHARD III and AMI, no collar tolerance is used. Note that our system automatically outputs SAD information, which however may not be perfect. To ensure a fair comparison with diarization methods [9, 23, 19, 20] that utilize the oracle SAD, we apply the SAD post-processing method presented in [13] to the diarization results of the proposed model.
@@ V-B1 Results on CALLHOME
Table IV shows the results on CALLHOME. The proposed model achieves new SOTA performance compared to other online methods. The performance advantages get more clear along with the increase of speakers. Considering that other online methods all use Sim{1-4}spk at the pre-training stage, we also provide the results (in grey) of the proposed model adapted from the pre-trained Sim{1-4}spk model for reference, which also show certain performance advantages.
@@ V-B2 Results on DIHARD II and DIHARD III
Table V and VI show the results on DIHARD II and DIHARD III, respectively. These two datasets are challenging in the sense that up to 9 speakers can be involved in the test utterances. The proposed model achieves the best performance among online methods under most conditions, which again verifies that the proposed model can simply handle a large number speakers by progressively increasing the training speakers. For the case with oracle SAD on DIHARD III, the proposed model performs better than EEND-GLA-Small and Conformer OTS-VAD, but comparable or worse than their variants, i.e. EEND-GLA-Large and ResNet OTS-VAD. The encoder architecture of the proposed model, i.e. four stacks of Conformer blocks, is more similar to the one of EEND-GLA-Small and Conformer OTS-VAD. These results show that leveraging a larger encoder as in EEND-GLA-Large or a CNN encoder as in ResNet OTS-VAD is possibly more suitable for some scenarios.
@@
In the proposed training strategy, model adaption with real-world datasets is always based on the short-utterance pre-trained model. One alternative model for real-world adaption could be the one already fine-tuned with simulated long utterances, e.g. the last model in Table III. The results of this setting (long pre-train) are also provided in Table V and VI, where adaption is performed directly using 400 s real-world data. It can be seen that this alternative performs slightly worse than the proposed one, which means it is better to conduct model adaption starting with short utterances.
@@ V-B3 Results on AMI
AMI has an average utterance duration of 33 minutes, which pose significant challenges especially for online diarization systems to maintain the speaker consistency over such a long time period. As discussed in Section III-D, the block-wise online diarization methods [26, 25] can simply process long audio streams block by block, however the difficulty of resolving the speaker permutation ambiguity between chunks gets higher for them with the increase of audio length. This work aims to leverage one single model to process long audio streams frame by frame, and speaker embeddings and attractors are kept consistent along the entire stream. This is achieved by training with long utterances in a progressive way. To analyze the effectiveness of this training strategy, DER of the proposed model w.r.t audio length with different training strategies are plotted in Fig. 3. Note that the DER curves were smoothed using the Savizky-Golay filter [50] for visualization purpose. It can be seen that the 200s model suffers from the length extrapolation problem, namely the performance decreases when test audio length is larger than training audio length. Directly fine-tuning with 1600 s data solves the length extrapolation problem, but degrades the performance of short audio part. The proposed progressive training strategy, i.e. 200s-$>$400s-$>$800s-$>$1600s, gradually improve the performance of longer audio without degrading the performance of short audio part. This demonstrates that there does exist an optimization point being optimal for both the short and long audio parts, and the proposed training strategy is effective for reaching such optimal point. When the progress step is set to a larger value, the 400s-$>$1600s model has a prominent performance decrease. Table VII shows the comparison results on AMI. We can see that the proposed model outperforms other online systems by a large margin.
Figure 3: DER curve with respect to audio length, for the AMI Dev set. ‘-$>$’ denotes the next step of adaptation.
@@
V-C Computational Efficiency Analysis
Additionally, we compare RTFs with several SOTA online methods in Table VIII. The cascaded clustering-based methods [20, 19] usually exhibit a high RTF. [19] improves the computational efficiency by utilizing restored checkpoint state in agglomerative hierarchy clustering (AHC). In block-wise online diarization methods [26, 49], normally a large block length is required to ensure the processing accuracy, while a small block shift is required to have a small processing latency. As a result, due to the large block overlap, the computational efficiency becomes lower, as every frame needs to be processed block lengthblock $\frac{\text{block length}}{\text{block shift}}$ times. In contrast, the proposed model processes once per frame, resulting in the RTF about an order of magnitude lower than that of EEND-EDA + FW-STB when using the same computing device, and several times lower than other methods although some of them use GPUs or multiple threads.
@@ V-C Computational Efficiency Analysis
EEND-EDA + FW-STB [25] AMD EPYC 7742 64-Core CPU @ 2.25 GHz 1 0.247
@@ V-D Ablation Study
Table IX shows the results of some ablation experiments conducted on Sim{1-4}spk data. When evaluated with appearance order, the DERs get higher. This performance degradation reflects the enrollment failure rate for speakers appearing, with the failure rate rising as the number of speaker increases. The effectiveness of L2-normalization, embedding similarity loss, look-ahead and convolution module in Conformer are validated, as the DERs get higher when any of these components is removed. To verify our new setting of $\gamma=1$ in Retention, its performance is compared with the original $\gamma$ setting as specified in (3). The significant performance degradation on 4 speakers when setting $\gamma$=Eq.(3) indicates that information decay over time is detrimental to long-speech diarization.
Furthermore, Fig. 5 visualizes and compares the speaker embeddings for different ablation cases. It can be seen that each component enhances the embedding distribution with better clustering of same speakers and discrimination between different speakers. Specifically, the embedding similarity loss improves clustering for individual speaker and also causes embeddings of overlapping speech to be distributed in between single-speaker embeddings. The look-ahead mechanism helps to remove some outliers. When $\gamma$=Eq.(3), the embeddings for same speakers exhibit noticeable discontinuities, indicating that information decay can impair the model’s ability to maintain speaker consistency over long periods of time.
@@
Methods Number of speakers
1 2 3 4
LS-EEND (prop.) 0.45 4.14 7.99 11.34
with appearance order 0.45 5.68 11.39 17.46
w/o L2-normalization 0.45 4.10 8.17 11.86
w/o embedding similarity loss 0.45 4.74 10.69 15.16
w/o look-ahead 0.49 5.48 10.67 14.03
w/o convolution module 0.46 4.94 11.87 16.77
with $\gamma$=Eq.(3) 0.37 3.92 8.95 18.74

### v2

@@ IV-A2 Real-world Recordings
TABLE II: Real-world Recordings. Recording (Rec) duration is presented in the format of minimum/average/maximum duration. Ovl stands for overlap ratio.
Dataset Split #Spk #Rec Ovl (%) Duration (mins)
CALLHOME [52] Part 1 2-7 249 17.0 0.86/2.10/10.11
Part 2 2-6 250 16.7 0.77/2.05/10.01
DIHARD II [53] Dev 1-10 192 9.8 0.45/7.44/11.62
Test 1-9 194 8.9 0.63/6.96/13.50
DIHARD III [54] Dev 1-10 254 10.7 0.45/8.07/11.62
Test 1-9 259 9.4 0.63/7.65/13.50
AMI headset mix [55] Train 3-5 136 13.4 7.97/35.59/90.25
@@ IV-B Model Configuration
Following the feature extraction method presented in [12, 27], the input audio features are set to 23-dimensional LogMels for every 10 ms, which are then concatenated with left and right 7 frames, and downsampled by a factor of 10, resulting in a 345-dimensional feature for every 0.1 s. The number of stacked blocks in encoder and decoder are set to NN=4 and MM=2, respectively. For both the encoder and decoder, the number of attention heads and hidden units are set to 4 and 256, respectively. The feed-forward dimension is set to 1024 in the encoder and 2048 in the decoder. The convolution kernel size in the Conformer module is set to 16, with a zero-padding of 15 to ensure causality. In the look-ahead module, the kernel size and padding are set to 19 and 9, respectively. Consequently, each frame attends to its previous and future 9 frames. Overall, the processing latency is $(1+9)\times 0.1+7\times 0.01$ = 1.07 s. In the following, we approximate the overall delay as 1 s for presentation simplicity. As already mentioned, the decay factor $\gamma$ in Retention is set to 1.
@@ IV-C Training Setup
The training process is divided into two stages: short-recording pre-training and long-segment adaptation. Pre-training is progressively conducted with linearly increased number of speakers, using the very large simulated dataset, and the length of training segments is set to 100 s. Specifically, the model is trained in multiple steps with the subsets of Sim2spk, Sim{1-4}spk, Sim{1-6}spk and Sim{1-8}spk for 100, 50, 50 and 25 epochs, respectively. Training segments are randomly extracted from the original recordings in an on-the-fly manner. For each step, the Noam optimizer [47] with 100,000 warm-up steps is utilized.
The pre-trained model cannot directly generalize to longer audio streams and real-world data. At the adaption stage, the pre-trained models are fine-tuned and adapted to specific applications/datasets. The fine-tuning process is also progressively conducted, but now with doubly increased audio lengths, using respective adaption data of each dataset. For real-world datasets, adaption is completed simultaneously for longer audio streams and real-world data. Audio length is increased from 200 s until it approximates the average recording duration of each dataset. Specifically, audio length is progressively set to 200 s and 400 s for the simulated dataset, to 200 s for CALLHOME, to 200 s and 400 s for DIHARD II and DIHARD III, to 200 s, 400 s, 800 s, and 1600 s for AMI. The adaption segments are also randomly extracted from the original recordings in an on-the-fly manner. When training with long sequence, the chunk-wise recurrent paradigm described in Section III-B is used for the Retention modules, with a recurrent chunk length of 50 s. The Adam optimizer [56] with a learning rate of $1\times 10^{-5}$ is used. In our experiments, we find that the model fine-tuned using the BCE loss (according to the speaker appearance order) does not converge well when adapted to real-world recordings. Therefore, we utilize the PIT loss [15] for real-world datasets.
As mentioned in Section III-E, during the inference stage, the model outputs activity probabilities corresponding to the maximum number of attractors. To enable the network automatically handle a flexible number of speakers below this upper bound, the termination marker (as illustrated in Fig. 1) is adopted. During training, for a training sample with $s$ ($s\leq S$) actual speakers, the (s+1)(s{+}1)-th attractor is designated as the termination marker (with all zero labels), and the diarization loss is computed on the 0-th to the (s+1)(s{+}1)-th attractor. By training the network to be aware of an explicit termination marker, then the network can automatically count active speakers at inference.
@@ IV-D Evaluation Setup
Diarization error rate (DER) is adopted as the evaluation metric, which is computed after performing the optimal speaker matching between the prediction and ground-truth label. The decision of diarization results is made using a threshold of 0.5. For the simulated data and CALLHOME, a collar tolerance of 0.25 s is used. For DIHARD II, DIHARD III and AMI, no collar tolerance is used. Note that our system automatically outputs SAD information, which however may not be perfect. To ensure a fair comparison with diarization methods [9, 23, 19, 20] that utilize the oracle SAD, we apply the SAD post-processing method presented in [13] to the diarization results of the proposed model.
@@ V Experimental Results
This section presents the results. The fully trained model, namely pre-training to 8 speakers and fine-tuning with long audio, is considered as our final proposal, referred to as LS-EEND (prop.). We also provide comparisons with some other configurations of the proposed model, which will be referred to as LS-EEND (*), where * specifies the configuration. For all comparison methods, the results are directly quoted from their papers or their subsequent papers by the same authors. Except that our previously proposed FS-EEND model [31] is re-trained using data presented in this work.
@@ V-B1 Results on CALLHOME
Table IV shows the results on CALLHOME. The proposed model achieves new SOTA performance compared to other online methods. The performance advantages get more clear along with the increase of speakers. Considering that other online methods all used Sim{1-4}spk at the pre-training stage, we also provide the results (in grey) of the proposed model adapted from the pre-trained Sim{1-4}spk model for reference, which also show certain performance advantages.
@@ V-B2 Results on DIHARD II and DIHARD III
Table V and VI show the results on DIHARD II and DIHARD III, respectively. These two datasets are challenging in the sense that up to 9 speakers can be involved in the test recordings. The proposed model achieves the best performance among online methods under most conditions, which again verifies that the proposed model can simply handle a large number speakers by progressively increasing the training speakers. For the case with oracle SAD on DIHARD III, the proposed model performs better than EEND-GLA-Small and Conformer OTS-VAD, but comparable or worse than their variants, i.e. EEND-GLA-Large and ResNet OTS-VAD. The encoder architecture of the proposed model, i.e. four stacks of Conformer blocks, is more similar to the one of EEND-GLA-Small and Conformer OTS-VAD. These results show that leveraging a larger encoder as in EEND-GLA-Large or a CNN encoder as in ResNet OTS-VAD is possibly more suitable for some scenarios.
@@
In the proposed training strategy, model adaption with real-world datasets is always based on the short-segment pre-trained model. One alternative model for real-world adaption could be the one already fine-tuned with simulated long segments, e.g. the last model in Table III. The results of this setting (long pre-train) are also provided in Table V and VI, where adaption is performed directly using 400 s real-world data. It can be seen that this alternative performs slightly worse than the proposed one, which means it is better to conduct model adaption starting with short segments.
@@ V-B3 Results on AMI
AMI has an average recording duration of 33 minutes, which pose significant challenges especially for online diarization systems to maintain the speaker consistency over such a long time period. As discussed in Section III-D, the block-wise online diarization methods [27, 26] can simply process long audio streams block by block, however the difficulty of resolving the speaker permutation ambiguity between chunks gets higher for them with the increase of audio length. This work aims to leverage one single model to process long audio streams frame by frame, and speaker embeddings and attractors are kept consistent along the entire stream. This is achieved by training with long segments in a progressive way. To analyze the effectiveness of this training strategy, DER of the proposed model w.r.t audio length with different training strategies is plotted in Fig. 3. Note that the DER curves were smoothed using the Savizky-Golay filter [57] for visualization purpose. It can be seen that the 200s model suffers from the length extrapolation problem, namely the performance decreases when test audio length is larger than training audio length. Directly fine-tuning with 1600 s data solves the length extrapolation problem, but degrades the performance of short audio part. The proposed progressive training strategy, i.e. 200s->>400s->>800s->>1600s, gradually improve the performance of longer audio without degrading the performance of short audio part. This demonstrates that there does exist an optimization point being optimal for both the short and long audio parts, and the proposed training strategy is effective for reaching such optimal point. When the progress step is set to a larger value, the 400s->>1600s model has a prominent performance decrease. Table VII shows the comparison results on AMI. We can see that the proposed model outperforms other online systems by a large margin.
Figure 3: DER curve with respect to audio length, for the AMI Dev set. ‘->>’ denotes the next step of adaptation.
@@
EEND-EDA + FW-STB [26]* AMD EPYC 7742 64-Core CPU @ 2.25 GHz 1 0.247
@@
  • 2
[…]
V-C Computational Efficiency Analysis
Additionally, we compare RTFs with several SOTA online methods in Table VIII. The cascaded clustering-based methods [20, 19] usually exhibit a high RTF. [19] improves the computational efficiency by utilizing restored checkpoint state in agglomerative hierarchy clustering (AHC). In block-wise online diarization methods [27, 28], normally a large block length is required to ensure the processing accuracy, while a small block shift is required to have a small processing latency. As a result, due to the large block overlap, the computational efficiency becomes lower, as every frame needs to be processed $\frac{\text{block length}}{\text{block shift}}$ times. In contrast, the proposed model processes once per frame, resulting in the RTF about an order of magnitude lower than that of EEND-EDA + FW-STB when using the same computing device, and several times lower than other methods although some of them use GPUs or multiple threads.
@@ V-D Ablation Study
Methods Simulated datasets Real-word datasets
Diarization loss Number of speakers Diarization loss CALLHOME DIHARD II DIHARD III AMI
1 2 3 4 All ($\leq$ 4) $\leq$ 4 $\leq$ 4 $\leq$ 4 Dev Eval
LS-EEND (Sim{1-4}spk) BCE 0.45 4.14 7.99 11.34 7.59 PIT 11.67 21.56 14.56 23.63 22.39
BCE 14.07 24.48 16.45 40.53 36.08
PIT 0.77 6.89 13.40 18.79 12.63 PIT 14.32 25.38 16.98 33.74 29.07
evaluation with appearance order BCE 0.45 5.68 11.39 17.46 11.25 - - - - - -
w/o L2-normalization BCE 0.45 4.10 8.17 11.86 7.65 PIT 11.71 21.73 14.62 23.24 23.59
w/o embedding similarity loss BCE 0.45 4.74 10.69 15.16 9.97 PIT 12.77 23.89 16.70 27.54 27.57
w/o look-ahead BCE 0.49 5.48 10.67 14.03 9.68 PIT 12.76 22.97 15.74 29.07 27.71
w/o convolution module BCE 0.46 4.94 11.87 16.77 10.97 PIT 13.26 22.91 16.42 28.73 29.32
w/o non-speech marker BCE 0.47 3.90 8.63 11.86 7.93 PIT 12.15 21.80 15.23 23.60 23.39
w/o termination marker BCE 0.84 10.24 14.02 11.89 10.81 PIT 12.65 23.17 15.39 25.43 24.53
with $\gamma$=Eq.(5) BCE 0.37 3.92 8.95 18.74 10.68 PIT 17.61 29.57 23.88 41.63 45.37
with masked self-attention BCE 0.53 3.76 7.95 9.86 6.93 PIT 11.69 21.02 14.03 20.59 22.69
Figure 5: An example of predicted speaker activities (for two ground truth speakers) when using or not using the speaker termination marker.
Table IX shows the results of ablation experiments to verify the effectiveness of the various schemes proposed in this work. Due to the largely increased training cost with more speakers, all ablation studies are conducted on data with no more than 4 speakers. Specifically, the ablation models are pre-trained with up to Sim{1-4}spk data, and then adapted with real-world data containing no more than 4 speakers.
We first compare when training the model with BCE loss (according to the speaker appearance order) or PIT loss. Note that the diarization performance is always measured with the best matching speaker order. We can see that BCE noticeably outperforms PIT for pre-training with simulated datasets. However, when adapting the BCE pre-trained model to real-world datasets, PIT noticeably outperforms BCE. Due to the worse performance of PIT pre-training, PIT pre-training plus PIT adaption does not perform as well as BCE pre-training plus PIT adaption. Appearance order is a natural arrangement of speaker order for online diarization, which is also consistent with the causal learning of the diarization network, in the sense that the network can sequentially arrange speaker based on historical information for both speaker embedding learning and attractor learning. Note that, in EEND, the changing/local speaker embeddings are learned based on certain rules, which means the order of multiple speaker embeddings also does matter. However, generally speaking, BCE training is more challenging than PIT training, as BCE forces a certain constraint compared to PIT. When training with a large amount of simulated data, both BCE and PIT can be well trained to have a small generalization error from training to test. In addition, BCE achieves better speaker embeddings than PIT, as we have observed considerably smaller embedding similarity loss for BCE than for PIT (the embedding similarity losses are not shown here). As a result, BCE noticeably outperforms PIT on simulated test sets. By contrast, when adapting the model with a small amount of real-world data, the more challenging BCE task cannot be well trained, as it has a larger generalization error (which is not shown here) than PIT. Overall, using BCE with a large amount of simulated data for pre-training and using PIT for adaption (if only a small amount of real-world data are available) is the optimal training configuration for the proposed model.
Next, on top of the LS-EEND (Sim{1-4}spk) model with the optimal training configuration, i.e. BCE pre-training plus PIT adaption, we conduct ablation experiments by either removing one component (w/o one component) or replacing one component with an alternative (with an alternative). When evaluated with appearance order, the DERs get higher. This performance degradation reflects the enrollment failure rate for speakers appearing, with the failure rate rising as the number of speaker increases. The effectiveness of L2-normalization, embedding similarity loss, look-ahead, convolution module in Conformer, non-speech marker and termination marker are validated, as the DERs get higher (for most conditions) for both simulated and real-world data when any of these components is removed. When removing the embedding similarity loss, the performance degradation on simulated data becomes more prominent as the number of speakers increases, which indicates that it is indeed more difficult for distinguishing more speakers, and the proposed embedding similarity loss helps to mitigate this problem. The performance degradation is also significant for the longest AMI data, which indicates that the embedding similarity loss is also helpful for maintaining the consistence of speaker embeddings along a very long period. The considerable performance degradation caused by removing the non-speech marker suggests that the non-speech attractor helps improve the discriminability of speech and non-speech frames to some extent. Removing the termination marker causes a significant performance degradation when the number of active speakers is smaller than the maximum number (4 in this experiment), such as the results of 2 and 3 speakers in the simulated test sets. Fig. 5 shows an example where fake speakers are detected when there is no termination marker. In contrast, the detection of active speakers can be correctly terminated when the termination marker is used.
To verify our new setting of $\gamma=1$ in Retention, its performance is compared with the original $\gamma$ setting as specified in (5). The significant performance degradation on simulated data with 4 speakers and on real-world data when setting $\gamma$=Eq.(5) indicates that information decay over time is detrimental to long-speech diarization. When we replace the Retention module with masked self-attention, performance can be slightly improved, but with a quadratic temporal complexity. Note that the audio length when fine-tuning on AMI is increased only up to 800 s, due to the limit of GPU memory.
Figure 6: DER curve with respect to audio length of LS-EEND (Sim{1-4}spk)-w/o embedding similarity loss, for the AMI Dev set. ‘->>’ denotes the next step of adaptation.
In Fig. 6, we evaluate the effectiveness of the proposed progressive training strategy on one of the baseline systems, LS-EEND(Sim{1-4})-w/o embedding similarity loss. It is demonstrated that the proposed training strategy remains effective when applied to alternative models.
Furthermore, Fig. 7 visualizes and compares the speaker embeddings for different ablation cases. It can be seen that each component enhances the embedding distribution with better clustering of same speakers and discrimination between different speakers. Specifically, PIT (pre-)training achieves less concentrated intra-speaker embeddings and less distinct embeddings between single-speaker and overlapped speakers. The embedding similarity loss improves clustering for individual speaker and also causes embeddings of overlapping speech to be distributed in between single-speaker embeddings. The look-ahead mechanism helps to remove some outliers. When $\gamma$=Eq.(5), the embeddings for same speakers exhibit noticeable discontinuities, indicating that information decay can impair the model’s ability to maintain speaker consistency over long periods of time.